During the past 8 hour shift, a patient drinks 3/4 orange juice, 0.5L water, and 3 oz. ice cream bar.              She has continuous IV infusing at the rate of 50 ml/hr. She has a urinary output of 660 ml, 120 diarrhea, and has had 2T of nemesis. what is the patient 8 hour I and O? record using whole numbers for both intake and output.
Answer #1

Ans) Intake:

Oral: Orange juice - 75 ml

Water= 0.5 L = 500 ml

Ice cream = 3 oz = 3 × 30 = 90 ml

Oral intake= 75 + 500 + 90 = 665 ml

IV: 50 ml/hr = 50 × 8 = 400 ml

Total intake = 665 ml + 400 ml = 1065 ml

Output :

U/O = 660 ml

Diarrhea= 120 ml

Vomit = 2 T

Total output = 660ml + 120 ml = 780 ml + 2T vomit

.